OpenStudy (anonymous):

will give medals to anyone who tries! Find the altitude of an isosceles triangle with a vertex angle of 70° and a base of 246. Write answer to nearest tenth.

OpenStudy (anonymous):

here is the correct answer for you tan(35) = 123/x x = 123tan(35) x = 123 * 0.7 = 86.1

OpenStudy (anonymous):

Since the altitude of isosceles triangle will divide both vertex angle and base in half => Altitude = half base/ tan 35° = 175.7